# A monolithic first–order BSSNOK formulation of the Einstein–Euler equations and its solution with path-conservative finite difference CWENO schemes

### v1

@@ VI Acknowledgments
This work was financially supported by i) the Italian Ministry of Education, University and Research (MIUR) in the framework of the PRIN 2022 project High order structure-preserving semi-implicit schemes for hyperbolic equations and via the Departments of Excellence Initiative 2018–2027 attributed to DICAM of the University of Trento (grant L. 232/2016) and ii) by the Departments of Excellence Initiative 2022–2027 attributed to the Department of Mathematics of the University of Roma La Sapienza (grant L. 232/2016).
M.D. and G.P. are members of the INdAM GNCS group in Italy.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high–performance computational resources and support.
We would kindly like to thank Ilya Peshkov, Luciano Rezzolla and Federico Guercilena for inspiring discussions. The authors declare that not a single line of this paper has been written with AI. The Fortran source code of the computer software which has generated the results shown in this paper can be obtained under CC BY-NC 4.0 license from the authors after written request.

### v2

@@ VI Acknowledgments
This work was financially supported by i) the Italian Ministry of Education, University and Research (MIUR) in the framework of the PRIN 2022 project High order structure-preserving semi-implicit schemes for hyperbolic equations and via the Departments of Excellence Initiative 2018–2027 attributed to DICAM of the University of Trento (grant L. 232/2016) and ii) by the Departments of Excellence Initiative 2022–2027 attributed to the Department of Mathematics of the University of Roma La Sapienza (grant L. 232/2016). MD was also funded by the Fondazione Caritro via the project SOPHOS and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, Grant agreement No. ERC-ADG-2021-101052956-BEYOND.
The authors of this work are all members of the INdAM GNCS group in Italy.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high–performance computing (HPC) resources and technical support.
We would like to thank Ilya Peshkov, Luciano Rezzolla, Federico Guercilena and Han Zhang for inspiring discussions. The authors would also kindly like to thank the anonymous referee for the numerous constructive comments and questions, which contributed to improve the overall quality and clarity of this paper substantially. The authors declare that not a single line of this paper has been written with AI. The Fortran source code of the computer software which has generated the results shown in this paper can be obtained under CC BY-NC 4.0 license from the authors after written request.